OpenStudy (anonymous):

Jimmy just turned 50 and has retirement savings in an IRA. How many years will he have to wait to be able to withdraw money without penalty? A. 70.5 B. 20.5 C. 9.5 D. 59.5

OpenStudy (anonymous):

u can get this crap all online if u type it in Google... just saying

OpenStudy (anonymous):

lol "D

OpenStudy (tkhunny):

50 + 70.5 = 120.5 That can't be it. 50 + 20.5 = 70.5 That's probably a mandatory withdrawal or annuitization date. 50 + 9.5 = 59.5 Hmmm... Suspicious. 50 + 59.5 = 109.5 That can't be it.

OpenStudy (tkhunny):

...or, I can just spew it all night! :-) Actually, it's past my bedtime. You're on your own.